In an R-L-C series circuit a maximum current of 0.5 A is obtained by varying the value of inductance L. The supply voltage is fixed at 230 V, 50 Hz. When maximum current flows through the circuit, the voltage measured across the capacitor is 350 V. What are the values of the cricuit parameters ?

1
Expert's answer
2022-03-21T12:41:58-0400

The circuit is in resonance state

Resistance of the circuit is active resistance. Ohm law for it


"R=\frac{U}I"

From the resonance state we find capacity and inductance


"C=\frac{I}{\omega U}""L=\frac{U}{I\omega}"

Calculating

"R=\frac{230}{0.5}=460"

"C=\frac{0.5}{350\cdot 314}=4.55\cdot10^{-6}"

"L=\frac{350}{0.5\cdot 314}=2.23"

Answer R=460 "\Omega" C=4.55 "\mu F" L= 2.23 H